OpenStudy (anonymous):

if you help me i'll love you forever omg For problems 1 and 2, graph the data points. Draw the line of best fit and, using complete sentences, explain how you found it in 3-4 sentences. Write the equation for the line of best fit in standard form. (10, 150) (20, 275) (30, 340) (40, 400) (50, 475) Title: Intersection Cars Horizontal axis: Minutes elapsed Vertical axis: Total cars passing through okay i'm done with everything but how to Write the equation for the line of best fit in standard form. so please help:)

OpenStudy (merchandize):

If it is total cars passing through, these are not independent data points. You will have to take the difference between successive data points. i.e. 150, 125, 65, 60, 75 to get number of cars per ten minutes. Then your best fit line might behave like an ordinary least squares line. f(x)=(150−x)2+(125−x)2+(65−x)2+(60−x)2+(75−x)2 f(x)=1502+1252+652+602+752−2(150+125+65+60+75)x+5x2 f′(x)=−2(150+125+65+60+75)+10x=0 x=(150+125+65+60+75)/5 Then the line is TotalCars=Minutes∗(150+125+65+60+75)/50

OpenStudy (anonymous):

woahhh thats alot to take in :( to be honest i'm sorta lost now.. TotalCars=Minutes∗(150+125+65+60+75)/50 is written in standard form?

OpenStudy (merchandize):

yes...is it fine??

OpenStudy (anonymous):

i have no clue

OpenStudy (merchandize):

where you have lost??

OpenStudy (anonymous):

i think this is how i'm suppose to do the standard form

OpenStudy (anonymous):

i tried solving it earlier but then i got confused

OpenStudy (anonymous):

anybodyy??

OpenStudy (anonymous):

-_-